OpenStudy (anonymous):

lim as x approaches positive infinity (1-3/x)^x i get log2 but my book say the answer is e^-3

OpenStudy (anonymous):

how do you get that?

OpenStudy (zarkon):

\[\lim_{n\to\infty}\left(1+\frac{r}{n}\right)^{nt}=e^{rt}\] if you let r=-3 and t=1 you get the result

OpenStudy (anonymous):

what happen to the x exponent

OpenStudy (zarkon):

x is a dummy variable. call it what you want still the same result \[\lim_{x\to\infty}\left(1+\frac{r}{x}\right)^{xt}=e^{rt}\]

OpenStudy (anonymous):

no i understand that im talking about the original eqution how do you know which is which. like how do you get that from the orginal

OpenStudy (anonymous):

The limit cwrw238 and Zarkon are using is the precise definition of the number \(e\approx2.718\).

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

this would help https://answers.yahoo.com/question/index?qid=20100123142609AATKWxy

OpenStudy (anonymous):

just follow the steps :)

OpenStudy (freckles):

@eggshell You requested my help. Do you still need help?

OpenStudy (anonymous):

@eggshell yeah do you :) ?

OpenStudy (anonymous):

yes im looking at the link that was posted and im confused on what happens to the natural log and how it just jumps to -1-3 over x^2. its the picture o

OpenStudy (freckles):

so you know e^(ln(y))=y

OpenStudy (freckles):

we are going to ignore the e part until later

OpenStudy (freckles):

lets just look at \[\lim_{x \rightarrow \infty}\ln[(1-\frac{3}{x})^x]\]

OpenStudy (freckles):

so you know by property of log we can bring down that power in the inside there

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

explained here : http://www.wolframalpha.com/input/?i=%28%281-3%29%2Fx%29%5Ex

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}x \ln(1-\frac{3}{x})=\lim_{x \rightarrow \infty}\frac{\ln(1-\frac{3}{x})}{\frac{1}{x}}\] here we have 0/0

OpenStudy (freckles):

we need to use l'hosptal's rule

OpenStudy (freckles):

first what is the derivative of ln(1-3/x)?

OpenStudy (freckles):

and then what is the derivative of 1/x?

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \frac{(\ln(1-\frac{3}{x}))' }{(\frac{1}{x})'}=\]

OpenStudy (anonymous):

1/(1-3/x) and x?

OpenStudy (freckles):

\[(\ln(1-\frac{3}{x}))'=\frac{(1-\frac{3}{x})'}{1-\frac{3}{x}}=\frac{0+\frac{3}{x^2}}{1-\frac{3}{x}}=\frac{\frac{3}{x^2}}{1-\frac{3}{x}}=\frac{3}{x^2-3x}\]

OpenStudy (freckles):

the derivative of 1/x is not x

OpenStudy (freckles):

\[(\frac{1}{x})'=(x^{-1})'=-1x^{-1-1}=-1x^{-2}=\frac{-1}{x^2}\]

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \frac{(\ln(1-\frac{3}{x}))' }{(\frac{1}{x})'}=\lim_{x \rightarrow \infty}\frac{3}{x^2-3x} \cdot (-x^2)\]

OpenStudy (freckles):

see if you can find the limit now

OpenStudy (anonymous):

i get zero

OpenStudy (xapproachesinfinity):

No do it again

OpenStudy (xapproachesinfinity):

you have -3x^2/(x^2-3x) it is not zero

OpenStudy (anonymous):

to he limit I plug in zero that's all I get

OpenStudy (freckles):

why plug in 0?

OpenStudy (freckles):

also if you plug in 0 we get 0/0 but we shouldn't be pluggin in 0 anyways

OpenStudy (freckles):

we should know from earlier in calculus that 1/x^n approaches 0 as x approach infinity for n greater than or equal to 1

OpenStudy (anonymous):

why not I thought that's what we did with infinity

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}\frac{-3x^2}{x^2-3x} \ \lim_{x \rightarrow \infty}\frac{-3}{1-\frac{1}{x}}=\frac{\lim_{x \rightarrow \infty}(-3)}{\lim_{x \rightarrow \infty}(1)-\lim_{x \rightarrow \infty}( \frac{1}{x})}\]

OpenStudy (freckles):

though you didn't have to do that either.. you could have used l'hosptal on the -3x^2/(x^2-3x) thing

OpenStudy (freckles):

since you have pm infinity/ininifty

OpenStudy (freckles):

before doing that you could also cancel a factor of x

OpenStudy (freckles):

and still use l'hospital lol

OpenStudy (freckles):

but yeah you don't plug in 0 when you have x approaches something that isn't 0

OpenStudy (anonymous):

so I just cancel out? my teacher told me to plus ito plug in zero with infintiy

OpenStudy (freckles):

cancel out?

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}\frac{-3x^2}{x^2-3x} \ \lim_{x \rightarrow \infty}\frac{-3}{1-\frac{1}{x}}=\frac{\lim_{x \rightarrow \infty}(-3)}{\lim_{x \rightarrow \infty}(1)-\lim_{x \rightarrow \infty}( \frac{1}{x})} \\ =\frac{-3}{1-0}=\frac{-3}{1}=-3\]